one number is equal to two times a second number. two time the first number plus two times the second number is 24. if you let x stand for the first number and y for the second, what are the two number.
Answer by rmromero(383)


one number is equal to two times a second number. two time the first number plus two times the second number is 24. if you let x stand for the first number and y for the second, what are the two number.
What is asked in the problem?
What are the two numbers.
Given:
one number is equal to two times the second number.
two times the first number plus two times the second number is 24

Representation
Let x = the first number
y = the second number
Equation:
x = 2y
2x + 2y = 24
We can use substitution to solve for this.
2x + 2y = 24 divide all terms by 2 to simplify
x + y = 12
2y + y = 12
3y = 12
y = 4

x = 2y, y = 4
x = 2(4)
x = 8
Checking:
2x + 2y = 24
2(8) + 2(4) = 24
16 + 8 = 24
24 = 24 ------> True

Therefore:
first number = 8
second number = 4


Visit my website:
http://onlinemathtutorial.tripod.com
Thanks